# SustainDiffusion: Optimising the Social and Environmental Sustainability of Stable Diffusion Models

### v1

@@ 3.2. Search Operators
We employ NSGA2 as the search algorithm of SustainDiffusion (Deb, K., 2002). NSGA2 is a very popular global search algorithm and has been shown to be widely effective and efficient for many multi-objective optimisation problems (Harman et al., 2008). It follows the general behaviour of genetic evolutionary algorithms, where an initial population progressively evolves following Darwin’s theory of biological evolution. First, a population of randomly generated individuals is created. The quality of each individual is assessed using the fitness functions defined in Section 3.3. Next, individuals are sorted using a non-dominated sorting strategy, and the best individuals go to the evolution process. In particular, each individual goes through crossover and mutation operations with a given probability. The fitness of the newly generated individuals is again evaluated and the best ones move onto the next generation. The evolution process continues until a maximum number of generations is reached.
Regarding crossover, SustainDiffusion uses a Single Point Uniform Crossover operator with 80% probability. Given two individuals $I_{1}$ and $I_{2}$ with $n$ keys each, this operator randomly chooses an index $i$ and generates two new individuals $I_{1}^{\prime}$ and $I_{2}^{\prime}$ such that $I_{1}^{\prime}$ contains the $[0,i]$ key values of $I_{1}$ and the $[(i+1),n]$ key values from $I_{2}$, while the opposite holds for $I_{2}^{\prime}$. For mutation, SustainDiffusion uses a Random Mutation operator with 20% of probability. This operator replaces the key values of an individual with those of a newly generated individual. Each key has an inner probability of being replaced, which is also set to 20%. Finally, the selection rate of SustainDiffusion (i.e., the number of best individuals to select for the next generation) is set to 5. The selection rates, crossover, and mutation probabilities are based on previous studies that used search-based methods for SD models (Gong et al., 2024; Berger et al., 2023).
@@ 3.3. Fitness Functions
The fitness of each individual is evaluated by feeding the SD model with the hyperparameters and prompt defined in the individual and generating a set of images using this setting. In particular, following previous studies (d’Aloisio et al., 2024a; Fadahunsi et al., 2025), we generate 20 images for each individual. Next, we assess the quality and bias of the generated images as well as the energy consumed by SD to generate them. In the following, we discuss each fitness function in detail.
@@ 3.3.1. Image Quality
Automatically assessing the quality of artificially generated images is a challenging task (Mila-Ceron et al., 2025). In this study, we follow previous work (Berger et al., 2023; Gong et al., 2024) by assessing image quality using the YOLO object detection model (Ultralytics, 2023). Given an image, YOLO detects objects in that image with a given confidence level. Next, the image quality is computed as the average of the confidence level of each object detected. Formally, the image quality fitness function is defined as follows:
(1) image $\text{image quality}=\frac{1}{n}\sum_{i=1}^{n}C_{i}$
@@ 3.3.2. Bias Metrics
Assessing the gender and ethnicity bias of a text-to-image generation model can be done in two steps (Fadahunsi et al., 2025; Sami et al., 2023; Luccioni et al., 2023). First, given a set of images generated by a model, the gender and ethnicity of the people depicted in each image must be identified. Next, following a specific fairness definition, the gender and ethnicity bias can be computed using the identified labels.
In this work, the gender and ethnicity of a person depicted in an image are automatically identified using the BLIP Visual-Question-Answering model (Li et al., 2022), which has been proven to be effective for this kind of task (Fadahunsi et al., 2025; Luccioni et al., 2023). BLIP works by taking an image and a question related to that image as input and providing a single-word label as the answer to the question. Following previous studies (Fadahunsi et al., 2025; Sami et al., 2023; Luccioni et al., 2023), we employ a binary gender classification and an ethnicity classification based on the 2021 England and Wales Census. [^2] https://www.ethnicity-facts-figures.service.gov.uk/style-guide/ethnic-groups/ Although this gender and ethnicity classification partially reflects reality, we argue that detecting non-binary genders and multiple ethnicities in artificially generated images is more challenging and prone to errors (Fadahunsi et al., 2025). Following the same studies (Fadahunsi et al., 2025; Luccioni et al., 2023), we feed BLIP with the following prompt to detect gender:“Is the person in this image a Male or a Female?”, and with the following prompt to detect ethnicity: “Is the person in this image Arab, Asian, Black, or White?”.
This labelling process is performed for each of the 20 images generated by SD for a given individual. Next, the gender and ethnicity biases can be assessed. Following the same previous studies (Fadahunsi et al., 2025; Sami et al., 2023; Luccioni et al., 2023), we use the Statistical Parity (SP) definition of fairness to compute gender and ethnic bias in the images generated by SD. SP states that a system is fair if it provides an equal distribution of all possible classification labels across all individuals, despite their belonging to specific groups (d’Aloisio et al., 2023; Mehrabi et al., 2022).
@@
(2) gender $\text{gender bias}=|P_{male}-P_{female}|$
@@
(3) ethnic $\text{ethnic bias}=|max(P_{e})-min(P_{e^{\prime}})|$
@@ 3.3.3. Energy Consumption Metrics
Building on previous studies (Roque et al., 2024; Georgiou et al., 2022), we evaluated the energy consumption of SD by measuring both CPU and GPU energy usage and the time taken to generate a single image. In particular, like for Image Quality, the energy fitnesses for an individual are defined as the median CPU, GPU, and time consumption across the 20 images generated.
However, incorporating all three metrics as additional fitness functions in our algorithm would result in a total of six different fitness functions (image quality, gender bias, ethnic bias, and three energy metrics). This increase in complexity would make it more difficult to address the search problem and identify optimal solutions (Chen and Li, 2023). Therefore, we conducted a preliminary empirical investigation to determine whether one of the energy consumption metrics could serve as a proxy for the others. In particular, we performed three different single-objective experiments by running SustainDiffusion each time using one of the three energy metrics as a fitness function. Since this search problem is single objective, we use Tournament Selection (Yang and Soh, 1997) as the selector operator with the number of selected individuals equal to five. The other search operators and settings are the same as reported in Section 3.2. For each experimental setting, we ran SustainDiffusion ten times and collected the CPU, GPU and time consumption as well as the image quality of each returned solution. After running all settings, we computed the Pareto-optimal set of all solutions based on the four metrics collected. The Pareto-optimal set is the set of solutions that are not dominated by any other solution in the search space, i.e., they are better in at least one objective and no worse in all other objectives (Harman et al., 2008). Next, we counted the number of times a solution returned by a search strategy appeared in the Pareto-optimal set.
@@ 3.4. Implementation Details
SustainDiffusion has been developed in Python 3.10 using only open-source tools and libraries. The search algorithm has been implemented using the deap Python library (Fortin et al., 2012). The SD version used in our experiments is Stable Diffusion 3 Medium from the Hugging Face repository, which is the latest version of SD available at the time of this study (Stability AI, 2024). The image quality has been assessed using YOLO v8 also from the Hugging Face repository (Ultralytics, 2023) and the gender and ethnic bias have been assessed using BLIP VQA Base still from Hugging Face (Li et al., 2022). Finally, the energy consumption has been measured using the Codecarbon Python library (Schmidt et al., 2020), which internally relies on pynvml and pyrapl to measure the GPU and CPU energy consumption, respectively.
@@ 4.1. Evaluation Process
To address RQ1, we compare the Pareto Optimal results obtained by SustainDiffusion with the results obtained by the original SD3 model from the Hugging Face repository with the default hyperparameters, and the results obtained by a Random Search (RS) algorithm that, at each iteration, selects a random individual from the search space and evaluates its fitness. Following research standards (Harman et al., 2008), the number of individuals evaluated by the RS algorithm at each iteration is equal to the average number of evaluations performed by SustainDiffusion among all generations, which, in our case, was equal to four. To tackle the non-stochastic behaviour of the algorithms, each approach has been executed 10 times. We recall that Pareto-optimal solutions returned by SustainDiffusion are those that are not dominated by any other solution in the search space - i.e., they are better than all other solutions in at least one fitness score and no worse in the others (Harman et al., 2008).
@@ 4.2. Evaluation Metrics
RQ1 and RQ2 are addressed by analysing both the individual objectives and the trade-offs between them. In addition to the objectives that SustainDiffusion directly optimises as fitness functions, we also consider GPU Energy and Time Duration to provide a more comprehensive overview of energy consumption. Individual objectives are analysed using descriptive statistics - i.e., mean and standard deviation - as well as distribution analysis using boxplots and statistical significance tests. For the statistical significance test, we employ the non-parametric Wilcoxon Sign-Ranked test to compare the averages of each objective across ten runs. The Wilcoxon test is a non-parametric test that makes no assumptions about the underlying data distribution (Wilcoxon, 1992). Therefore, it is a better solution for relatively small samples where the normal distribution is more difficult to achieve (Arcuri and Briand, 2011). The null hypothesis that we assess is: $H_{0}:$ The objective $O$ obtained by SustainDiffusion is not improved with respect to the baseline approach $x$. The alternative hypothesis is: $H_{1}:$ The objective $O$ obtained by SustainDiffusion is improved with respect to the baseline approach $x$. For Image Quality, ”improved” means that the score obtained by SustainDiffusion is higher than the baseline approach. For the other objectives, ”improved” means that the score obtained by SustainDiffusion is lower. Following standards (Sarro et al., 2016; Hort et al., 2024; Harman et al., 2008), we set the confidence level to $0.05$ and apply the Bonferroni correction to mitigate the risk of Type I statistical error – i.e., incorrectly rejecting the null hypothesis. Thus, we consider the test significant if its $p-$value is $<0.05/6$, where 6 is the number of evaluations performed for each SustainDiffusion-baseline pair (i.e., one for each objective). Finally, when the test is significant, we complement it with the non-parametric Vargha-Delaney $\hat{A}_{12}$ effect size to assess the magnitude of the difference (Vargha and Delaney, 2000). Following previous works (Sarro et al., 2016; Hort et al., 2024), we consider the following classes for the effect size: $\hat{A}_{12}\geq 0.72$: large, $0.720.64\leq\hat{A}_{12}<$ 0.72: medium, and $\hat{A}_{12}<0.64$: small.
@@
The trade-off between objectives is instead assessed using Pareto-optimality and the Hypervolume score. Pareto-optimality is computed by identifying the Pareto Front from all the solutions returned by all algorithms and counting the number of times a solution from a given algorithm $x$ appears in the front (Harman et al., 2008). Hypervolume is a metric used in multi-objective optimisation problems to evaluate the quality of Pareto-optimal solutions in the objective space with respect to a given reference point (Guerreiro et al., 2021). Higher values indicate better coverage of the objective space. The steps performed to compute the Hypervolume score are the following. First, for every run, we identify the Pareto Front from the solutions obtained by each approach. Thus, for each approach evaluated, we have ten Pareto Fronts (one for each run). Next, we aggregate all the Pareto Fronts from all approaches and identify the reference point by selecting the worst value of each objective plus an $\epsilon$ value of $0.5$. Finally, for each run, we compute the Hypervolume score using all Pareto Fronts and the reference point. We plot the distribution of the scores and compute the Wilcoxon test to assess if the Hypervolume achieved by SustainDiffusion is statistically significantly larger than the baselines.
@@
To address RQ3, we conduct the Kruskal-Wallis H test to determine if there are statistically significant differences in the scores obtained by SustainDiffusion over the ten runs. The Kruskal-Wallis H test is a non-parametric test used to verify the null hypothesis that the population medians of multiple ($\geq 2$) groups are equal (Kruskal and Wallis, 1952). When the test reported a statistically significant difference ($p-$value $<0.05$), we computed the post-hoc Dunn’s test to count the number of pairs reporting a statistically significant difference. Dunn’s test is a post-hoc test commonly used in conjunction with the Kruskal-Wallis H test to perform pairwise comparisons of mean ranks (Dunn, 1964). As done in the previous experiments, we applied the Bonferroni correction to mitigate Type I errors.
@@
Finally, to evaluate RQ4, following previous studies (Hort et al., 2024; Sarro et al., 2016), we conduct the so-called Win-Tie-Loss analysis. For each prompt and baseline approach, we count the number of times SustainDiffusion provides scores that dominate the baseline (Win), dominate the baseline in at least half the objectives (Tie), or are dominated by the baseline (Loss). We report the Win-Tie-Loss for each SustainDiffusion-baseline pair as well as the total number of wins, ties and losses obtained by SustainDiffusion.
@@ 4.3. Evaluation Data
The set of prompts used in our evaluation is the Software Engineering Tasks dataset from (Fadahunsi et al., 2025; Sami et al., 2023). This dataset contains a list of 56 different prompts of the form: ”Photo portrait of a Software Engineer that $<$task$>$, where $<$task$>$ represents a software-related activity, such as ”Fixes bugs”, ”Reviews pull requests”, or ”Edits code” among others. This dataset has been shown to expose significant gender and ethnic bias in SD models, especially through Male and White/Asian figures (Fadahunsi et al., 2025). In particular, the prompt used in RQ1, RQ2 and RQ3 is ”Photo portrait of a Software Engineer that codes”, while for RQ4, we employed the full set of prompts.

### v2

@@ 3.2. Search Operators
We employ NSGA2 as the search algorithm of SustainDiffusion (deb2002nsga2). NSGA2 is a very popular global search algorithm and has been shown to be widely effective and efficient for many multi-objective optimisation problems (harman2008search). It follows the general behaviour of genetic evolutionary algorithms, where an initial population progressively evolves following Darwin’s theory of biological evolution. First, a population of randomly generated individuals is created. The quality of each individual is assessed using the fitness functions defined in Section 3.3. Next, individuals are sorted using a non-dominated sorting strategy, and the best individuals go to the evolution process. In particular, each individual goes through crossover and mutation operations with a given probability. The fitness of the newly generated individuals is again evaluated and the best ones move onto the next generation. The evolution process continues until a maximum number of generations is reached.
Regarding crossover, SustainDiffusion uses a Single Point Uniform Crossover operator with 80% probability. Given two individuals $I_{1}$ and $I_{2}$ with $n$ keys each, this operator randomly chooses an index $i$ and generates two new individuals $I_{1}^{\prime}$ and $I_{2}^{\prime}$ such that $I_{1}^{\prime}$ contains the $[0,i]$ key values of $I_{1}$ and the $[(i+1),n]$ key values from $I_{2}$, while the opposite holds for $I_{2}^{\prime}$. For mutation, SustainDiffusion uses a Random Mutation operator with 20% of probability. This operator replaces the key values of an individual with those of a newly generated individual. Each key has an inner probability of being replaced, which is also set to 20%. Finally, the selection rate of SustainDiffusion (i.e., the number of best individuals to select for the next generation) is set to 5. The selection rates, crossover, and mutation probabilities are based on previous studies that used search-based methods for SD models (gong2024greenstableyolooptimizinginferencetime; berger2023stableyolo).
@@ 3.3. Fitness Functions
The fitness of each individual is evaluated by feeding the SD model with the hyperparameters and prompt defined in the individual and generating a set of images using this setting. In particular, following previous studies (d2024exploring; fadahunsi_how_2025), we generate 20 images for each individual. Next, we assess the quality and bias of the generated images as well as the energy consumed by SD to generate them. In the following, we discuss each fitness function in detail.
@@ 3.3.1. Image Quality
Automatically assessing the quality of artificially generated images is a challenging task (mila-ceron_easy_2025). In this study, we follow previous work (berger2023stableyolo; gong2024greenstableyolooptimizinginferencetime) by assessing image quality using the YOLO object detection model (yolov8). Given an image, YOLO detects objects in that image with a given confidence level. Next, the image quality is computed as the average of the confidence level of each object detected. Formally, the image quality fitness function is defined as follows:
(1) $\text{image quality}=\frac{1}{n}\sum_{i=1}^{n}C_{i}$
@@ 3.3.2. Bias Metrics
Assessing the gender and ethnicity bias of a text-to-image generation model can be done in two steps (fadahunsi_how_2025; sami_case_2023; luccioni_stable_2023). First, given a set of images generated by a model, the gender and ethnicity of the people depicted in each image must be identified. Next, following a specific fairness definition, the gender and ethnicity bias can be computed using the identified labels.
In this work, the gender and ethnicity of a person depicted in an image are automatically identified using the BLIP Visual-Question-Answering model (blip), which has been proven to be effective for this kind of task (fadahunsi_how_2025; luccioni_stable_2023). BLIP works by taking an image and a question related to that image as input and providing a single-word label as the answer to the question. Following previous studies (fadahunsi_how_2025; sami_case_2023; luccioni_stable_2023), we employ a binary gender classification and an ethnicity classification based on the 2021 England and Wales Census. [^2] https://www.ethnicity-facts-figures.service.gov.uk/style-guide/ethnic-groups/ Although this gender and ethnicity classification partially reflects reality, we argue that detecting non-binary genders and multiple ethnicities in artificially generated images is more challenging and prone to errors (fadahunsi_how_2025). Following the same studies (fadahunsi_how_2025; luccioni_stable_2023), we feed BLIP with the following prompt to detect gender:“Is the person in this image a Male or a Female?”, and with the following prompt to detect ethnicity: “Is the person in this image Arab, Asian, Black, or White?”.
This labelling process is performed for each of the 20 images generated by SD for a given individual. Next, the gender and ethnicity biases can be assessed. Following the same previous studies (fadahunsi_how_2025; sami_case_2023; luccioni_stable_2023), we use the Statistical Parity (SP) definition of fairness to compute gender and ethnic bias in the images generated by SD. SP states that a system is fair if it provides an equal distribution of all possible classification labels across all individuals, despite their belonging to specific groups (daloisio_debiaser_2023; mehrabi2022surveybiasfairnessmachine).
@@
(2) $\text{gender bias}=|P_{male}-P_{female}|$
@@
(3) $\text{ethnic bias}=|max(P_{e})-min(P_{e^{\prime}})|$
@@ 3.3.3. Energy Consumption Metrics
Building on previous studies (roque2024unveiling; georgiou2022green), we evaluated the energy consumption of SD by measuring both CPU and GPU energy usage and the time taken to generate a single image. In particular, like for Image Quality, the energy fitnesses for an individual are defined as the median CPU, GPU, and time consumption across the 20 images generated.
However, incorporating all three metrics as additional fitness functions in our algorithm would result in a total of six different fitness functions (image quality, gender bias, ethnic bias, and three energy metrics). This increase in complexity would make it more difficult to address the search problem and identify optimal solutions (10.1145/3514233). Therefore, we conducted a preliminary empirical investigation to determine whether one of the energy consumption metrics could serve as a proxy for the others. In particular, we performed three different single-objective experiments by running SustainDiffusion each time using one of the three energy metrics as a fitness function. Since this search problem is single objective, we use Tournament Selection (yang1997structural) as the selector operator with the number of selected individuals equal to five. The other search operators and settings are the same as reported in Section 3.2. For each experimental setting, we ran SustainDiffusion ten times and collected the CPU, GPU and time consumption as well as the image quality of each returned solution. After running all settings, we computed the Pareto-optimal set of all solutions based on the four metrics collected. The Pareto-optimal set is the set of solutions that are not dominated by any other solution in the search space, i.e., they are better in at least one objective and no worse in all other objectives (harman2008search). Next, we counted the number of times a solution returned by a search strategy appeared in the Pareto-optimal set.
@@ 3.4. Implementation Details
SustainDiffusion has been developed in Python 3.10 using only open-source tools and libraries. The search algorithm has been implemented using the deap Python library (DEAP_JMLR2012). The SD version used in our experiments is Stable Diffusion 3 Medium from the Hugging Face repository, which is the latest version of SD available at the time of this study (sd3). The image quality has been assessed using YOLO v8 also from the Hugging Face repository (yolov8) and the gender and ethnic bias have been assessed using BLIP VQA Base still from Hugging Face (blip). Finally, the energy consumption has been measured using the Codecarbon Python library (codecarbon), which internally relies on pynvml and pyrapl to measure the GPU and CPU energy consumption, respectively.
@@ 4.1. Evaluation Process
To address RQ1, we compare the Pareto Optimal results obtained by SustainDiffusion with the results obtained by the original SD3 model from the Hugging Face repository with the default hyperparameters, and the results obtained by a Random Search (RS) algorithm that, at each iteration, selects a random individual from the search space and evaluates its fitness. Following research standards (harman2008search), the number of individuals evaluated by the RS algorithm at each iteration is equal to the average number of evaluations performed by SustainDiffusion among all generations, which, in our case, was equal to four. To tackle the non-stochastic behaviour of the algorithms, each approach has been executed 10 times. We recall that Pareto-optimal solutions returned by SustainDiffusion are those that are not dominated by any other solution in the search space - i.e., they are better than all other solutions in at least one fitness score and no worse in the others (harman2008search).
@@ 4.2. Evaluation Metrics
RQ1 and RQ2 are addressed by analysing both the individual objectives and the trade-offs between them. In addition to the objectives that SustainDiffusion directly optimises as fitness functions, we also consider GPU Energy and Time Duration to provide a more comprehensive overview of energy consumption. Individual objectives are analysed using descriptive statistics - i.e., mean and standard deviation - as well as distribution analysis using boxplots and statistical significance tests. For the statistical significance test, we employ the non-parametric Wilcoxon Sign-Ranked test to compare the averages of each objective across ten runs. The Wilcoxon test is a non-parametric test that makes no assumptions about the underlying data distribution (wilcoxon1992individual). Therefore, it is a better solution for relatively small samples where the normal distribution is more difficult to achieve (arcuri_practical_2011). The null hypothesis that we assess is: $H_{0}:$ The objective $O$ obtained by SustainDiffusion is not improved with respect to the baseline approach $x$. The alternative hypothesis is: $H_{1}:$ The objective $O$ obtained by SustainDiffusion is improved with respect to the baseline approach $x$. For Image Quality, ”improved” means that the score obtained by SustainDiffusion is higher than the baseline approach. For the other objectives, ”improved” means that the score obtained by SustainDiffusion is lower. Following standards (sarro_multi-objective_2016; HortEMSE; harman2008search), we set the confidence level to $0.05$ and apply the Bonferroni correction to mitigate the risk of Type I statistical error – i.e., incorrectly rejecting the null hypothesis. Thus, we consider the test significant if its $p$-value is <0.05/6<0.05/6, where 6 is the number of evaluations performed for each SustainDiffusion-baseline pair (i.e., one for each objective). Finally, when the test is significant, we complement it with the non-parametric Vargha-Delaney $\hat{A}_{12}$ effect size to assess the magnitude of the difference (vargha2000critique). Following previous works (sarro_multi-objective_2016; HortEMSE), we consider the following classes for the effect size: $\hat{A}_{12}\geq 0.72$: large, $0.64\leq\hat{A}_{12}<0.72$: medium, and $\hat{A}_{12}<0.64$: small.
@@
The trade-off between objectives is instead assessed using Pareto-optimality and the Hypervolume score. Pareto-optimality is computed by identifying the Pareto Front from all the solutions returned by all algorithms and counting the number of times a solution from a given algorithm $x$ appears in the front (harman2008search). Hypervolume is a metric used in multi-objective optimisation problems to evaluate the quality of Pareto-optimal solutions in the objective space with respect to a given reference point (10.1145/3453474). Higher values indicate better coverage of the objective space. The steps performed to compute the Hypervolume score are the following. First, for every run, we identify the Pareto Front from the solutions obtained by each approach. Thus, for each approach evaluated, we have ten Pareto Fronts (one for each run). Next, we aggregate all the Pareto Fronts from all approaches and identify the reference point by selecting the worst value of each objective plus an $\epsilon$ value of $0.5$. Finally, for each run, we compute the Hypervolume score using all Pareto Fronts and the reference point. We plot the distribution of the scores and compute the Wilcoxon test to assess if the Hypervolume achieved by SustainDiffusion is statistically significantly larger than the baselines.
@@
To address RQ3, we conduct the Kruskal-Wallis H test to determine if there are statistically significant differences in the scores obtained by SustainDiffusion over the ten runs. The Kruskal-Wallis H test is a non-parametric test used to verify the null hypothesis that the population medians of multiple ($\geq 2$) groups are equal (kruskal1952use). When the test reported a statistically significant difference ($p$-value <0.05<0.05), we computed the post-hoc Dunn’s test to count the number of pairs reporting a statistically significant difference. Dunn’s test is a post-hoc test commonly used in conjunction with the Kruskal-Wallis H test to perform pairwise comparisons of mean ranks (dunn1964multiple). As done in the previous experiments, we applied the Bonferroni correction to mitigate Type I errors.
@@
Finally, to evaluate RQ4, following previous studies (HortEMSE; sarro_multi-objective_2016), we conduct the so-called Win-Tie-Loss analysis. For each prompt and baseline approach, we count the number of times SustainDiffusion provides scores that dominate the baseline (Win), dominate the baseline in at least half the objectives (Tie), or are dominated by the baseline (Loss). We report the Win-Tie-Loss for each SustainDiffusion-baseline pair as well as the total number of wins, ties and losses obtained by SustainDiffusion.
@@ 4.3. Evaluation Data
The set of prompts used in our evaluation is the Software Engineering Tasks dataset from (DALOISIO2026107956; fadahunsi_how_2025; sami_case_2023). This dataset contains a list of 56 different prompts of the form: ”Photo portrait of a Software Engineer that <<task>>, where <<task>> represents a software-related activity, such as ”Fixes bugs”, ”Reviews pull requests”, or ”Edits code” among others. This dataset has been shown to expose significant gender and ethnic bias in SD models, especially through Male and White/Asian figures (DALOISIO2026107956; fadahunsi_how_2025). In particular, the prompt used in RQ1, RQ2 and RQ3 is ”Photo portrait of a Software Engineer that codes”, while for RQ4, we employed the full set of prompts.